Show that, when SI units for μ0 and ϵ0 are entered, the units given by the right-hand side of the equation in the problem above are m/s.


1
Expert's answer
2022-01-24T11:18:47-0500

We know that

"c=\frac{1}{\sqrt{\mu_0\epsilon_0}}\rightarrow(1)"

RHS

Unit of

"\mu_0=\frac{T\times m}{A}"

"\epsilon=\frac{c^2}{N\times m^2}"

"A=\frac{c}{sec}"

"\mu_0=\frac{T\times sec\times m}{c}"

Equation (1) put RHS value

"c=\frac{1}{\sqrt{\frac{T\times m \times sec}{c} \times{\frac {c^2}{N\times m^2}}}}"

"c=\frac{1}{\sqrt{\frac{T \times sec\times c}{N\times m}}}\rightarrow(2)"

Now We know that

"F=qvB"

"v=\frac{F}{qB}"

Unit

"v=\frac{N}{c\times T}\rightarrow({3})"

Equation (2) and (3) we can written as

"c=\frac{1}{\sqrt{\frac{1}{v}\times \frac{sec}{m}}}"

"\frac{1}{v}=\frac{sec}{m}"

"c=\frac{1}{\sqrt{\frac{1}{v^2}}}"

"c=v"

Unit of c=v =m/sec

LHS=RHS